OpenStudy (anonymous):

solve for x: 3(x + 1) = -2( x -1) + 6

OpenStudy (anonymous):

3x+3 = -2x+2 + 6 3x+3 = -2x+8 5x+3=8 5x = 5 x = 1

OpenStudy (anonymous):

that make sense?